hello, could you be so kind to help me solve these two problems explaining the steps since i do not know how to solve them please!
1. A train ticket in a certain city is $3.00. People who use the train also have the option of purchasing a frequent rider pass for $15.00 each month. With the pass, each ticket costs only $2.25. Determine the number of times in a month the train must be used so that the total monthly cost without the pass is the same as the total monthly cost with the pass.


2. After a 17% price reduction, a boat sold for $20,750. what was the boat's price before the reduction> (round to the nearest cent, if necessary)

Answer by advanced_Learner(501)

hello, could you be so kind to help me solve these two problems explaining the steps since i do not know how to solve them please!
1. A train ticket in a certain city is $3.00. People who use the train also have the option of purchasing a frequent rider pass for $15.00 each month. With the pass, each ticket costs only $2.25. Determine the number of times in a month the train must be used so that the total monthly cost without the pass is the same as the total monthly cost with the pass.

with out pass
ticket cost is 3x$
with the pass
cost is 15+2.25x$
15+2.25x=3
1500+225x=300x
1500=75x
x=20
must be used 20 times or more.


2. After a 17% price reduction, a boat sold for $20,750. what was the boat's price before the reduction> (round to the nearest cent, if necessary)

original price is
x-0.17x=20,750
x(1-0.17)=20,750
x(0.83)=20,750
x=25000